OpenStudy (anonymous):

FAN AND MEDAL!! Please help Solve x2 + 9x + 20 = 0. x = 2, x = 10 x = -2, x = -10 x = -5, x = -4 x = 5, x = 4

OpenStudy (jdoe0001):

hint: \(\large \begin{array}{ccclllll} x^2 &+ 9x &+ 20 &= 0\\ &\uparrow &\uparrow \\ &5+4&5\cdot 4 \end{array}\)

OpenStudy (anonymous):

So would it be D? @jdoe0001

OpenStudy (studygurl14):

Actually, it would be C. Because using the information @jdoe0001 gave, you'd do this: \(\large (x+5)(x+4)=0\) Then split into two equations and solve for x \(\large x+5 = 0\) and \(\large x+4=0\)

OpenStudy (studygurl14):

Which would result in \(\large x=-5,-4\)

OpenStudy (anonymous):

ohhh ok I see what I did wrong thanks @StudyGurl14 ...

OpenStudy (studygurl14):

You're welcome. :)